# for the Baryon Scattering (BaSc) Collaboration

### v1

@@ Abstract
We perform a high-statistics lattice QCD calculation of the low-energy two-nucleon scattering amplitudes. In order to address discrepancies in the literature, the calculation is performed at a heavy pion mass in the limit that the light quark masses are equal to the physical strange quark mass, $m_{\pi}=m_{K}\simeq 714$ MeV. Using a state-of-the-art momentum space method, we rule out the presence of a bound di-nucleon in both the isospin 0 (deuteron) and 1 (di-neutron) channels, in contrast with many previous results that made use of compact hexaquark creation operators. In order to diagnose the discrepancy, we add such hexaquark interpolating operators to our basis and find that they do not affect the determination of the two-nucleon finite volume spectrum, and thus they do not couple to deeply bound di-nucleons that are missed by the momentum-space operators. Further, we perform a high-statistics calculation of the HAL QCD potential on the same gauge ensembles and find qualitative agreement with our main results. We conclude that two-nucleons do not form bound states at heavy pion masses and that previous identification of deeply bound di-nucleons must have arisen from a misidentification of the spectrum from off-diagonal elements of a correlation function.
preprint: LLNL-JRNL-2005660
@@ I Introduction
To access the two-particle scattering amplitudes, the finite-volume formalism in the form of the two-body Quantization Condition (QC2) Luscher (1986, 1991) is used to relate the two-particle spectrum in finite volume to the infinite volume scattering amplitudes. Specifically, what is required is a precise and accurate determination of the interacting energy of the system. A significant challenge for two-nucleon (NN) calculations is that this interaction energy is $\mathrm{O}(0.1\%)$ of the total energy of the system. Because LQCD is a relativistic QFT, and calculations are carried out with quark and gluon degrees of freedom, we must compute the total energy of the system rather than having direct access to these small interacting energies. In and of itself, this is not a problem, but LQCD calculations involving nucleons also suffer from an exponentially bad signal-to-noise (S/N) problem Parisi (1984); Lepage (1989); in the large time limit, the S/N degrades roughly as ${\rm S/N}\sim\sqrt{N_{\rm sample}}\,{\rm exp}(-A(m_{N}-\frac{3}{2}m_{\pi})t)$ where $A$ is the number of nucleons.
@@
For small Euclidean times, correlation functions receive significant contributions from excited states. For single hadrons, the excited state gap is roughly set by $2m_{\pi}$. Thus, the ground state is resolvable over a Euclidean time of roughly $1-2$ fm. In contrast, for NN calculations, the excited state gap is roughly set by the quantized momentum modes of the nucleon. With periodic spatial boundary conditions, these are ${\Delta}E_{n}=2\sqrt{m_{N}^{2}+p_{n}^{2}}-2m_{N}\approx p_{n}^{2}/m_{N}$ with $p_{n}=\sqrt{n}(2\pi/L)$ for integer $n$. For typical sizes of the lattice $L$, these energy gaps are $\mathrm{O}(20-50)$ MeV and thus, without a suppression of the excited states, the ground state is not resolvable until $4-10$ fm in time while the S/N for NN calculations has degraded around $t\sim 2$ fm. Thus, for correlation functions constructed from a single creation/annihilation operator, there is no region in time where the ground state dominates the signal and the S/N is sufficiently large that the ground state interaction energy can be precisely or accurately determined.
@@
This S/N problem has contributed to discrepancies in the literature, even at heavy pion masses of ${m_{\pi}\approx 800}$ MeV, where the S/N challenge is significantly less severe. Notably, different groups report qualitatively conflicting results on whether di-nucleons form bound states. In particular, after the first dynamical LQCD calculation of the NN system was performed Beane et al. (2006), an alternative method was proposed to determine the NN interactions Ishii et al. (2007) which is known as the HAL QCD Potential Ishii et al. (2012); Aoki and Doi (2020). Earlier calculations of NN systems indicated deeply bound di-nucleons Beane et al. (2012, 2013a); Yamazaki et al. (2012, 2015); Berkowitz et al. (2017); Beane et al. (2017); Wagman et al. (2017) while results with the HAL QCD potential found no bound states Iritani et al. (2016, 2017, 2019a).
@@
This discrepancy was believed to be due to uncontrolled systematic uncertainties with the HAL QCD Potential method Detmold et al. (2007); Beane (2008); Beane et al. (2008, 2011); Walker-Loud (2014); Savage (2016). At the same time, subsequent work by HAL QCD raised serious concerns about the calculations that observed bound states Iritani et al. (2016, 2017); Aoki et al. (2018); Iritani et al. (2019a). This was followed up by two-baryon calculations that utilized a set of operators that enabled the use of momentum space creation operators, in contrast to the local hexa-quark (HX) creation operators used in Ref. Beane et al. (2012, 2013a); Yamazaki et al. (2012, 2015); Berkowitz et al. (2017); Beane et al. (2017); Wagman et al. (2017) (these operators will be defined subsequently). These new methods found significantly reduced binding energies in the case of the h-dibaryon Francis et al. (2019); Green et al. (2021), or a lack of bound states in the case of di-nucleons Hörz et al. (2021); Amarasinghe et al. (2023). This discrepancy has long plagued progress toward the physical point and led to a lack of confidence in results reported for further two- and higher- nucleon quantities Drischler et al. (2019); Tews et al. (2022).
@@ II NN spectrum and amplitudes
Scattering amplitudes are generally not directly accessible in LQCD calculations (though see Ref. Bulava and Hansen (2019) for a proposal to do so with smeared spectral functions). Instead, the two-particle finite volume quantization condition Luscher (1986, 1991) (QC2), often referred to as “the Lüscher method”, provides a mapping between the two-particle spectrum and the infinite volume scattering amplitudes: in essence, the two-particle interactions cause the energies to be shifted, in a finite-volume, from their corresponding non-interacting values providing a relationship between the scattering phase shifts and geometric functions that depend upon the spatial geometry and boundary conditions. Therefore, the essential ingredient in determining the scattering amplitudes with LQCD is the spectrum. In particular, the small shift of the two-particle energy from the non-interacting levels must be determined. For NN systems, ${\Delta}E_{NN}=E_{NN}-2m_{N}$ is $\sim\mathrm{O}(0.1\%)$ of the total energy, presenting significant challenges for an accurate and precise determination given the known excited state contamination and S/N degradation challenges in LQCD calculations.
@@
$\mathbf{d}_{\rm tot}^{2}$ irrep ${\rm N}_{op}$ irrep ${\rm N}_{op}$
@@ 4 $E$ 15
The energies of the stationary states in finite-volume are discrete and are determined using the standard Hermitian correlator matrix method Fox et al. (1982); Ishikawa et al. (1982); Michael and Teasdale (1983); Michael (1985); Luscher and Wolff (1990) in lattice QCD. Given a set of $N_{op}$ operators, an $N_{op}\times N_{op}$ matrix of two-point correlation functions is evaluated for a range of temporal separations $t$ using a Markov-chain Monte Carlo method. The elements of the Hermitian correlator matrix have spectral representations given by
@@
The Hermitian correlator matrix method described above is sometimes referred to as a “variational” method in the literature. We emphasize that the variational method, which is an approximation scheme in quantum mechanics to obtain an energy estimate by minimizing the matrix element of the Hamiltonian in a parametrized wave function, is not used for extracting the energies in the method above, so the systematic errors of the variational method do not apply. The pivoted-matrix above can be viewed as a change of operator basis to one which can be deemed optimal from a variational viewpoint. Euclidean time projection guarantees that, as long as the overlap onto the lowest lying state with the operator basis is not exactly zero, this state will be resolved for long enough times within the Hermitian correlator matrix method. Further, a larger operator basis will provide better control over excited state contamination than a limited operator basis.
@@ II.1.1 Lattice Action
This calculation uses the CLS lattice action Bruno et al. (2015), which is an $N_{f}=2+1$ isotropic clover-Wilson action that is non-perturbatively $\mathrm{O}(a)$ improved, and the Lüscher-Weisz improved gluon action with tree-level coefficients was used with periodic boundary conditions in all directions. The lattice spacing is $a\approx 0.086$ fm with a lattice volume of $48^{3}\times 96$. The three quark masses are all approximately equal to the physical strange quark mass, i.e. $m_{u}=m_{d}=m_{s}\approx m_{s}^{\rm phys}$, which leads to $m_{\pi}L\approx 15$. This ensemble was generated in Ref. Hörz et al. (2021) and named C103.111In the CLS naming scheme settled after they were generated, this ensemble would be named C150. A total of 1490 configurations are used from four replicas/streams. The choice of quark masses was made to roughly match the SU(3) symmetric calculation of NPLQCD Beane et al. (2013a) where it was claimed that deeply bound di-nucleons existed Beane et al. (2013a); Berkowitz et al. (2017); Wagman et al. (2017).
@@ II.1.2 Interpolating operator construction
When $m_{u}=m_{d}=m_{s}$ and QED interactions are ignored, QCD has an $SU(3)$ flavor symmetry. $SU(3)$ flavor basis states $|\Lambda,I,I_{3},Y\rangle$ can be labeled by irrep $\Lambda$ (the quadratic and cubic Casimirs), isospin $I$, isospin projection $I_{3}$, and hypercharge $Y$. Note that $Y=S+B$, where $S$ is strangeness and $B$ is baryon number. The baryons $N^{0}$, $N^{+}$, $\Lambda^{0}$, $\Sigma^{-}$, $\Sigma^{0}$, $\Sigma^{+}$, $\Xi^{-}$, $\Xi^{0}$ then all have the same mass since they all reside in a flavor $\bm{8}$ irrep. For $NN$ scattering at the $SU(3)$ symmetric point, we must consider the $\bm{8}\times\bm{8}$ flavor irrep, whose decomposition into irreps is
@@
The majority of our calculations in our other studies are carried out using $SU(2)$ flavor symmetry since we usually employ the more realistic approximation $m_{u}=m_{d}$ with $m_{s}$ being much heavier. A very significant amount of work has gone into designing the $SU(2)$ operators and into the software evaluating the correlations of such operators. Hence, we decided to utilize our $SU(2)$ operators even for this calculation at the $SU(3)$ symmetric point. The states produced by our $SU(2)$ nucleon operators should have substantial overlap with those produced by $SU(3)$ nucleon operators and should lead to the same lowest-lying energy in each isospin channel. Another way to view our use of $SU(2)$ nucleon operators is to note that we have not initially block-diagonalized the spectrum as much as could be done with $SU(3)$ operators, but the same spectrum should be obtained in the space spanned by the states created by the operators.
@@ II.2 Correlator Analysis: Conspiracy Model
In all previous LQCD calculations of two-baryons, one consistent observation is that the excited state contamination of the two-baryon correlation function is nearly the same as that of the single-baryons. For example, the effective mass of the ratio correlation function exhibits mild excited state contamination and often plateaus earlier in time than either the two or single baryon correlators. This has led to a common strategy of fitting the ratio correlation functions to extract the interaction energies, $\Delta E=E_{N_{1}N_{2}}-E_{N_{1}}-E_{N_{2}}$ Aoki et al. (2010); Berkowitz et al. (2017); Iritani et al. (2019a); Green et al. (2021); Hörz et al. (2021); Green et al. (2023). Such a strategy is criticized by the NPLQCD Collaboration Wagman et al. (2017); Amarasinghe et al. (2023) because, even if the two-nucleon correlator is positive-definite, and thus monotonically approaches the large time limit, the ratio correlation function is not guaranteed to do so.
@@
where both the numerator and denominator were described with a sum of exponentials with positive overlap factors, $r_{n}=A_{n}/A_{0}$ with $A_{n}=\langle n|O|{\Omega}\rangle$,
@@
$C_{NN}(t)=\left[\sum_{n_{1}=0}^{N_{1}-1}A_{n}^{(1)}e^{-E_{n}^{(1)}t}\right]\left[\sum_{n_{2}=0}^{N_{2}-1}A_{n}^{(2)}e^{-E_{n}^{(2)}t}\right]\,.$
@@
    $E^{N^{n}_{1}N^{m}_{2}}=E_{n}^{N_{1}}+E_{m}^{N_{2}}+\delta_{nm}\,,$ (11)
    where $E_{n}^{N_{i}}=E_{0}^{N_{i}}+\Delta_{n0}^{N_{i}}$ is the $n^{th}$ energy of nucleon $i$ and $\delta_{nm}$ is an interaction energy between the $n^{th}$ state of nucleon-1 ($N_{1}$) and the $m^{th}$ state of nucleon-2 ($N_{2}$).
@@
    A global analysis is performed to simultaneously constrain the single nucleon ground and excited state energies, $E_{n}^{N_{i}}$, along with the two-nucleon interacting energies, $\delta_{nm}$ as well as the overlap factors for all correlators in a given irrep.
@@
We contrast this conspiracy model with an agnostic model in which the number of excited states in the NN correlator are not fixed by the choice of the single nucleon correlators, nor are the energies coupled as in Eq. (11). Our previous work implemented such an agnostic model (and also fit the ratio correlation functions). In this work, we show that the conspiracy models are strongly favored over the agnostic models as measured by the Bayes Factor in a Bayesian analysis.

### v2

@@ Abstract
We perform a high-statistics lattice QCD calculation of the low-energy two-nucleon scattering amplitudes. In order to address discrepancies in the literature, the calculation is performed at a heavy pion mass in the limit that the light quark masses are equal to the physical strange quark mass, $m_{\pi}=m_{K}\simeq 714$ MeV. Using a state-of-the-art momentum space method, we rule out the presence of a bound di-nucleon in both the isospin 0 (deuteron) and 1 (di-neutron) channels, in contrast with many previous results that made use of compact hexaquark creation operators. In order to diagnose the discrepancy, we add such hexaquark interpolating operators to our basis and find that they do not affect the determination of the two-nucleon finite volume spectrum, and thus they do not couple to deeply bound di-nucleons that are missed by the momentum-space operators. Further, we perform a high-statistics calculation of the HAL QCD potential on the same gauge ensembles and find qualitative agreement with our main results. We conclude that di-nucleons do not form bound states at heavy pion masses and that previous identification of deeply bound di-nucleons must have arisen from a misidentification of the spectrum from off-diagonal elements of a correlation function.
@@ I Introduction
To access the two-particle scattering amplitudes, the finite-volume formalism in the form of the two-body Quantization Condition (QC2) Luscher (1986, 1991) is used to relate the two-particle spectrum in finite volume to the infinite volume scattering amplitudes. Specifically, what is required is a precise and accurate determination of the interacting energy of the system. A significant challenge for two-nucleon (NN) calculations is that this interaction energy is $\mathrm{O}(0.1\%)$ of the total energy of the system. Because LQCD is a relativistic quantum field theory (QFT), and calculations are carried out with quark and gluon degrees of freedom, we must compute the total energy of the system rather than having direct access to these small interacting energies. This is not a problem in principle, but LQCD calculations of temporal correlation functions involving nucleons also suffer from an exponentially bad signal-to-noise (S/N) problem Parisi (1984); Lepage (1989): in the large time limit, the S/N degrades roughly as ${\rm S/N}\sim\sqrt{N_{\rm sample}}\,{\rm exp}(-A(m_{N}-\frac{3}{2}m_{\pi})t)$ where $A$ is the number of nucleons.
@@
For small Euclidean times, correlation functions receive significant contributions from excited states. For single hadrons, the excited state gap is roughly set by $2m_{\pi}$. Thus, the ground state is resolvable over a Euclidean time of roughly 121-2 fm. In contrast, for NN calculations, the excited state gap is roughly set by the quantized momentum modes of the nucleon. With periodic spatial boundary conditions, these are ${\Delta}E_{n}=2\sqrt{m_{N}^{2}+p_{n}^{2}}-2m_{N}\approx p_{n}^{2}/m_{N}$ with $p_{n}=\sqrt{n}(2\pi/L)$ for integer $n$. For typical sizes of the lattice $L$, these energy gaps are $\mathrm{O}(20-50)$ MeV and thus, without a suppression of the excited states, the ground state is not resolvable until 4104-10 fm in time while the S/N for NN calculations has degraded around $t\sim 2$ fm. Thus, for correlation functions constructed from a single creation/annihilation operator, there is no region in time where the ground state dominates the signal and the S/N is sufficiently large that the ground state interaction energy can be precisely or accurately determined.
@@
This S/N problem has contributed to discrepancies in the literature, even at heavy pion masses of $m_{\pi}\approx 800$} MeV, where the S/N challenge is significantly less severe. Notably, different groups report qualitatively conflicting results on whether di-nucleons form bound states, for example, see the 2012 and 2013 reviews from the annual lattice conference Doi (2012); Walker-Loud (2014). In particular, after the first dynamical LQCD calculation of the NN system was performed Beane et al. (2006), an alternative method was proposed to determine the NN interactions Ishii et al. (2007) which is known as the HAL QCD Potential Ishii et al. (2012); Aoki and Doi (2020). Earlier calculations of NN systems indicated deeply bound di-nucleons Beane et al. (2012, 2013a); Yamazaki et al. (2012, 2015); Berkowitz et al. (2017); Beane et al. (2017); Wagman et al. (2017) while results with the HAL QCD potential found no bound states Iritani et al. (2016, 2017, 2019a).
@@
This discrepancy was believed to be due to uncontrolled systematic uncertainties with the HAL QCD Potential method Detmold et al. (2007); Beane (2008); Beane et al. (2008, 2011); Walker-Loud (2014); Savage (2016). At the same time, subsequent work by HAL QCD raised serious concerns about the calculations that observed bound states Iritani et al. (2016, 2017); Aoki et al. (2018); Iritani et al. (2019a). This was followed up by two-baryon calculations that utilized a set of operators that enabled the use of momentum space creation operators, in contrast to the local hexa-quark (HX) creation operators used in Ref. Beane et al. (2012, 2013a); Yamazaki et al. (2012, 2015); Berkowitz et al. (2017); Beane et al. (2017); Wagman et al. (2017) (these operators will be defined subsequently). These new methods found significantly reduced binding energies in the case of the $H$ dibaryon Francis et al. (2019); Green et al. (2021), and a lack of bound states in the case of di-nucleons Hörz et al. (2021); Amarasinghe et al. (2023). This discrepancy has long plagued progress toward the physical point and led to a lack of confidence in results reported for further two- and more-nucleon scattering amplitudes Drischler et al. (2021); Tews et al. (2022).
@@ II NN spectrum and amplitudes
Scattering amplitudes are generally not directly accessible in LQCD calculations (though see Refs. Bulava and Hansen (2019); Patella and Tantalo (2025) for a proposal to do so with smeared spectral functions). Instead, the two-particle finite volume quantization condition Luscher (1986, 1991) (QC2), often referred to as “the Lüscher method”, provides a mapping between the two-particle spectrum and the infinite volume scattering amplitudes: in essence, the two-particle interactions cause the energies to be shifted, in a finite-volume, from their corresponding non-interacting values providing a relationship between the scattering phase shifts and geometric functions that depend upon the spatial geometry and boundary conditions. Therefore, the essential ingredient in determining the scattering amplitudes with LQCD is the spectrum. In particular, the small shift of the two-particle energy from the non-interacting levels must be determined. For NN systems, ${\Delta}E_{NN}=E_{NN}-2m_{N}$ is $\sim\mathrm{O}(0.1\%)$ of the total energy, presenting significant challenges for an accurate and precise determination given the known excited state contamination and S/N degradation challenges in LQCD calculations.
@@
$\mathbf{d}_{\rm tot}^{2}$ irrep ${\rm N}_{\rm op}$ irrep ${\rm N}_{\rm op}$
@@ 4 $E$ 15
The energies of the stationary states in finite-volume are discrete and are determined using the standard Hermitian correlator matrix method Fox et al. (1982); Ishikawa et al. (1982); Michael and Teasdale (1983); Michael (1985); Luscher and Wolff (1990) in lattice QCD. Given a set of $N_{\rm op}$ operators, an $N_{\rm op}\times N_{\rm op}$ matrix of two-point correlation functions is evaluated for a range of temporal separations $t$ using a Markov-chain Monte Carlo method. The elements of the Hermitian correlator matrix have spectral representations given by
@@
The Hermitian correlator matrix method described above is sometimes referred to as a “variational” method in the literature. However, we emphasize that this should not be confused with the variational method in quantum mechanics, which is an approximation scheme used to obtain an energy bound by minimizing the expectation value of the Hamiltonian in a parametrized wave function. The variational method in quantum mechanics does not guarantee convergence to the lowest-lying energies of the system when using a fixed number of wave-function ansätze. By contrast, the pivoted matrix introduced above can be interpreted as a change of operator basis to one that is, in a sense, optimal from a variational perspective. In this case, however, Euclidean time projection ensures that, as long as the overlap of the operator basis with the lowest-lying state is non-zero, that state will be isolated at sufficiently large Euclidean time separations. Moreover, the functional form of the excited-state contamination is now explicitly determined by the Euclidean time dependence, and it is exponential with a large energy gap that depends on the size of the operator basis. Thus, the systematics of the two methods are quite different: the Hermitian correlator matrix method does not yield upper bounds, but rather estimates whose excited-state contamination can be systematically understood and suppressed via Euclidean time evolution. Furthermore, increasing the size of the operator basis improves control over this contamination by enlarging the energy gap that governs its suppression.
@@ II.1.1 Lattice Action
This calculation uses the CLS lattice action Bruno et al. (2015), which is an $N_{f}=2+1$ isotropic clover-Wilson action that is non-perturbatively $\mathrm{O}(a)$ improved, and the Lüscher-Weisz improved gluon action with tree-level coefficients was used with periodic boundary conditions in all directions. The lattice spacing is $a\approx 0.086$ fm with a lattice volume of $48^{3}\times 96$. The three quark masses are all approximately equal to the physical strange quark mass, i.e. $m_{u}=m_{d}=m_{s}\approx m_{s}^{\rm phys}$, which leads to $m_{\pi}L\approx 15$. This ensemble was generated in Ref. Hörz et al. (2021) and named C103.111In the CLS naming scheme settled after they were generated, this ensemble would be named C150. A total of 1490 configurations are used from four replica. The choice of quark masses was made to roughly match the SU(3) symmetric calculation of Ref. Beane et al. (2013a) where it was claimed that deeply bound di-nucleons existed Beane et al. (2013a); Berkowitz et al. (2017); Wagman et al. (2017).
@@ II.1.2 Interpolating operator construction
When $m_{u}=m_{d}=m_{s}$ and QED interactions are ignored, QCD has an $SU(3)$ flavor symmetry. $SU(3)$ flavor basis states $|\Lambda,I,I_{3},Y\rangle$ can be labeled by irrep $\Lambda$ (the quadratic and cubic Casimirs), isospin $I$, isospin projection $I_{3}$, and hypercharge $Y$. Note that $Y=S+B$, where $S$ is strangeness and $B$ is baryon number. The baryons $^{0}, N+N^{+}$, $\Lambda^{0}$, $\Sigma^{-}$, $\Sigma^{0}$, $\Sigma^{+}$, $\Xi^{-}$, $\Xi^{0}$ then all have the same mass since they all reside in a flavor $\bm{8}$ irrep. For $NN$ scattering at the $SU(3)$ symmetric point, we must consider the $\bm{8}\times\bm{8}$ flavor irrep, whose decomposition into irreps is
@@
The majority of our calculations in our other studies are carried out using $SU(2)$ flavor symmetry since we usually employ the more realistic approximation $m_{u}=m_{d}$ with $m_{s}$ being much heavier. A significant amount of work has gone into designing the $SU(2)$ operators and into the software evaluating the correlations of such operators. Hence, we utilize our $SU(2)$ operators even for this calculation at the $SU(3)$ symmetric point. The states produced by our $SU(2)$ nucleon operators should have substantial overlap with those produced by $SU(3)$ nucleon operators and should lead to the same lowest-lying energy in each isospin channel. Another way to view our use of $SU(2)$ nucleon operators is to note that we have not initially block-diagonalized the spectrum as much as could be done with $SU(3)$ operators, but the same spectrum should be obtained in the space spanned by the states created by the operators.
@@ II.2 Correlator Analysis: Conspiracy Model
In all previous LQCD calculations of two-baryons, one consistent observation is that the excited state contamination of the two-baryon correlation function is nearly the same as that of the single-baryons. For example, the effective mass of the ratio correlation function exhibits mild excited state contamination and often plateaus earlier in time than either the two or single baryon correlators. This has led to a common strategy of fitting the ratio correlation functions to extract the interaction energies, $\Delta E=E_{N_{1}N_{2}}-E_{N_{1}}-E_{N_{2}}$ Aoki et al. (2010); Berkowitz et al. (2017); Iritani et al. (2019a); Green et al. (2021); Hörz et al. (2021); Green et al. (2023). Such a strategy is criticized, for example in Refs. Wagman et al. (2017); Amarasinghe et al. (2023), because even if the two-nucleon correlator is positive-definite, and thus monotonically approaches the large time limit, the ratio correlation function is not guaranteed to do so.
@@
where both the numerator and denominator were described with a sum of exponentials with positive overlap factors, $r_{n}=z_{n}/z_{0}$ with $z_{n}=\langle n|O|{\Omega}\rangle$, $A_{n}=|z_{n}|^{2}$,
@@
$C_{NN}(t)=\left[\sum_{n_{1}=0}^{N_{1}-1}A_{n}^{(1)}e^{-E_{n}^{(1)}t}\right]\left[\sum_{n_{2}=0}^{N_{2}-1}A_{n}^{(2)}e^{-E_{n}^{(2)}t}\right]\,.$ (11)
@@
    $E^{N^{n}_{1}N^{m}_{2}}=E_{n}^{N_{1}}+E_{m}^{N_{2}}+\delta E_{nm}\,$, (12)
    where $E_{n}^{N_{i}}=E_{0}^{N_{i}}+\Delta_{n0}^{N_{i}}$ is the $n^{th}$ energy of nucleon $i$ and $\delta E_{nm}$ is an interaction energy between the $n^{th}$ state of nucleon-1 ($N_{1}$) and the $m^{th}$ state of nucleon-2 ($N_{2}$).
@@
    A global analysis is performed to simultaneously constrain the single nucleon ground and excited state energies, $E_{n}^{N_{i}}$, along with the two-nucleon interacting energies, $\delta E_{nm}$ as well as the overlap factors for all correlators in a given irrep.
@@
We contrast this conspiracy model with an agnostic model in which the number of excited states in the NN correlator are not fixed by the choice of the single nucleon correlators, nor are the energies coupled as in Eq. (12). Our previous work implemented such an agnostic model (and also fit the ratio correlation functions). In this work, we show that the conspiracy models are favored over the agnostic models as measured by the Bayes Factor in a Bayesian analysis.